Echelle Failure

Monday, June 20, 2016 2:26 PM

Problem(s) Encountered:

Telescope would not record data from the slit and only recorded zeros, not even a bias level. Called Eric Galayda at beginning of night to try to find a solution. he tried restarting the computer, restarting the electronics on the telescopes and none of these seemed to help. The electronics box was then shut down for an hour to let it cool down, but at this point KPNO had closed the 4m due to smoke so the dome and telescope remained off for the rest of

the night.


Turns out to be a problem with Echelle.  It looks like the problem is within the electronics box for the detector, which apparently is specific in use to Echelle.  Found nothing obvious upon inspection of all boards and components.  Reseated everything but the problem persists.  The electronics box has been sent to ARC and we are awaiting word on what they find out.